OpenStudy (anonymous):

What is the acceleration of a ball traveling horizontally with an initial velocity of 20 meters/second and, 2.0 seconds later, a velocity of 30 meters/second?

OpenStudy (anonymous):

I think the best way to approach these physics problems is to first write everything we know down assigned to variables we see in useful equations. In this case, we know that V1 (for initial velocity) = 20m/s V2 (final velocity) = 30 m/s and that the change in time is ∆t = 2s. Since you know two velocities in two points in time you can find the change in velocity ∆V = V2 - V1. So the change in velocity (∆V) = 10m/s. Knowing the definition of acceleration (a = ∆V ÷ ∆t) you know that the acceleration is 10÷2 and therefore = 5m/s/s. You can check your answer using units knowing that you divided meters per second by seconds. Your answer should then be meters per second per second.